OpenStudy (anonymous):

solve the equation for solutions over the interval [0 degrees,360 degrees] tan^2 theta+5tan theta+3=0

OpenStudy (freckles):

Can you solve u^2+5u+3=0?

OpenStudy (freckles):

\[\tan(\theta)=u=\frac{-5 \pm \sqrt{25 -4(1)(3)}}{2(1)}\]

OpenStudy (anonymous):

do I factor u^2+5u+3=0?